# Combinatorial Selection with Costly Information

### v1

@@ 5 Pandora’s Box with Partial Inspection (Minimization)
Formally, we consider an instance with $n$ partial inspection boxes (henceforth, PI-boxes) $\{\mathcal{B}_{i}\}_{i=1}^{n}$, where each box $\mathcal{B}_{i}=(\mathcal{D}_{i},c^{o}_{i},c^{p}_{i})$ is characterized by a distribution $\mathcal{D}_{i}$ over values; an opening cost $c^{o}_{i}>0$; and a peeking cost $c^{p}_{i}\in(0,c^{o}_{i})$ [^9] Since we can only select an opened box, if $c^{p}_{i}\geq c^{o}_{i}$, then we can safely exclude the peeking action and the box reduces to a classical Pandora’s Box.. Each PI-box $\mathcal{B}_{i}$ can be expressed as a Costly Information MDP with two possible actions, peeking and opening, where we can interpret the accept action after peeking as incurring a cost of $c^{o}$. An instance of PBPI corresponds to an instance of min-CICS over the corresponding MDPs. A pictorial representation of the different states of a box and the underlying MDP is shown in Figure 3.
Figure 3: Each PI-box $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$ is initially closed. In order to learn the value realization realization $X\sim\mathcal{D}$, the decision maker can either peek into the box (at a cost of $c^{p}$) or open it (at a cost of $c^{o}$). To accept the box, the decision maker must first open it and then pay its (now known) value $X$.
Our Results.
In Section 5.1 we will show that while there are simple instances of PBPI where all committing algorithms are sub-optimal up to a constant factor (Example 2), there exists a trivial committing algorithm that achieves a $\phi\approx 1.618$ approximation to the optimal adaptive policy (Lemma 5.2). In order to further improve on this guarantee, in Section 5.2 we will develop a $\sqrt{2}\approx 1.414$ local approximation guarantee for any PI-box $\mathcal{B}$ (Lemma 5.3); combined with the meta-algorithm of Corollary 4.2, this directly implies our main result for this section:
@@
There exists an efficient algorithm that achieves a $\sqrt{2}$-approximation to the optimal adaptive policy for any instance of matroid-PBPI.
5.1 Lower and Upper Bounds on the Adaptivity Gap of PBPI
In this section, we take a closer look at committing algorithms for PBPI and the approximation guarantees that they can achieve with respect to the optimal adaptive policy. Observe that a committing algorithm for PBPI needs to decide in advance (perhaps randomly) whether each box $\mathcal{B}_{i}$ will be directly opened or peeked into and then (potentially) opened. We begin by showing that that there are simple instances where all committing policies are sub-optimal up to a constant factor.
@@ 5.1 Lower and Upper Bounds on the Adaptivity Gap of PBPI
  • Since opening the second box is free, we can assume without loss that all policies start by opening it and observing its value, call it $y$. The optimal policy will simply open and accept the first box if $y=\infty$. But if $y=2$, it can peek into the first box and only open it if it contains a value of 00. The expected cost of this algorithm is $\frac{15}{8}$.
  • Now consider any policy that commits to opening or peeking into box $1$ before observing $y$. If it commits to opening the first box, it accepts the value of this box regardless of $y$, as box $1$ always has a value smaller than $y$. The expected cost of this policy is $2$. If it commits to peeking into the first box, then at $y=\infty$ it incurs a cost of $\frac{9}{4}$ due to having to pay the extra peeking cost; and at $y=2$ it incurs an expected cost of $\frac{7}{4}$ by opening the first box only if it contains a value of 00. It’s net cost is again $2$.
@@
Example 2 illustrates that in general, committing algorithms are worse than the optimal adaptive policy by a factor of at least $\frac{16}{15}$. On the positive side, it is easy to obtain a $2$-approximation via committing algorithms. In particular, consider the policy that always commits to peeking. This policy can mimic the optimal one as follows. Whenever the optimal algorithm peeks, so does this committing policy. Whenever the optimal algorithm opens without peeking, the committing policy peeks and then opens; on a box with costs $c^{p}_{i}$ and $c^{o}_{i}$, this policy pays $c^{p}_{i}+c^{o}_{i}$ or at most twice the amount $c^{o}_{i}$ paid by the optimal algorithm. In fact, we can further refine this argument by choosing the action we commit to more carefully: there always exists a simple commitment under which we can achieve a $\phi\approx 1.618$ approximation to the optimal adaptive policy. The proof is presented in Appendix C.
@@
Consider any instance $\mathbb{I}=(\mathcal{B}_{1},\cdots,\mathcal{B}_{n},\mathcal{F})$ of matroid-PBPI and partition the $n$ boxes into two sets
@@
and $P=[n]\setminus O$. The policy that commits to directly opening the boxes in $O$ and peeking before opening the boxes in $P$ achieves a $\phi$-approximation to the optimal adaptive policy.
We note that by using a global argument such as the one above, i.e. an argument where we charge each action of the committing algorithm directly to the optimal adaptive policy, one cannot improve on this upper bound of $\phi$ for the gap between committing and adaptive policies (for example, consider a setting where all the boxes have $c^{p}=1$ and $c^{o}=\phi$). In order to improve on this guarantee, we would need to leverage our knowledge of the value distributions. In the next section, we achieve this by establishing local approximation guarantees for PBPI.
@@ 5.2 Local Approximation Guarantees for PBPI
In this section, we prove Theorem 5.1 by providing a $\sqrt{2}$-local approximation for any PI-box, as stated below. Given a PI-box $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$, let $g^{p}$ denote the water filling (a.k.a. Gittins) index of the policy that commits to peeking. Equivalently, $g^{p}$ is the solution to the equation $c^{p}=\operatorname{E}_{X\sim\mathcal{D}}\left[(g^{p}-X-c^{o})^{+}\right]$. We call $g^{p}$ the peeking index of the box.
@@
Let $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$ be a PI box with peeking index $g^{p}$. Let $\pi$ be the policy that commits to the opening action whenever
@@
Definition 9.
The optimality curve of a PI-box $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$ for cost minimization is given by
@@
Figure 4: The optimality curves for a PI-box $\mathcal{B}$ with opening cost $c^{o}=0.5$, value $X=0$ with probability $0.5$ and $X=2$ with probability $0.5$ and peeking cost $c^{p}=0.4$ (left) and $c^{p}=0.1$ (right). The optimality curve of $\mathcal{B}$ is given by the minimum of the three curves. Observe that the curves of the opening and peeking action intersect at a unique point $\tau$ and that unless $g^{p}<g^{o}$, the opening action dominates the peeking action.
Observe that if the optimal adaptive policy prefers opening over peeking the box for some outside option $y$, then the same will be true for any $y^{\prime}>y$. Also, for $y=0$ peeking is clearly preferable to opening as $c^{p}<c^{o}$. Thus, the optimality curves $f^{o}_{\mathcal{B}}(y)$ and $f^{p}_{\mathcal{B}}(y)$ will have a unique intersection (up to an interval). We use $\tau$ to denote this intersection; formally, $\tau$ is the maximal solution to equation
This implies that unless $g^{p}<g^{o}<\tau$, the opening action will dominate the peeking action for all values of the outside option $y$ for which accepting it isn’t optimal; as a consequence, such instances trivially admit a $1$-local approximation by committing to the opening action. From now on, we assume that $g^{p}<g^{o}<\tau$. The, Lemma 5.3 is a consequence of the following claims, establishing local approximation guarantees for the opening and peeking actions respectively.
@@
Note that once proven, they immediately imply an $\alpha$-local approximation with
@@
Proof of 1.
By concavity, we have that $f^{o}_{\mathcal{B}}(y)\leq\alpha f^{o}_{\mathcal{B}}(\frac{y}{\alpha})$ for all $y\in\real$ and $\alpha\geq$ 1; thus, we only need to verify the condition for $y$ such that $f_{\mathcal{B}}(\frac{y}{\alpha})=f^{p}_{\mathcal{B}}(\frac{y}{\alpha})$. In other words, we only need to verify that $\min\{y,f^{o}_{\mathcal{B}}(y)\}\leq\alpha f^{p}_{\mathcal{B}}(\frac{y}{\alpha})$ for all $y\in[\alpha g^{p},\alpha\tau]$.
For start, we will show that the condition holds for $\alpha=g^{o}/g^{p}$; notice that for this parameter, $f^{o}_{\mathcal{B}}(y)\leq y$ for all $y\in[\alpha g^{p},\alpha\tau]$ and thus we can now write our condition as $D_{\alpha}(y):=\alpha f^{p}_{\mathcal{B}}(\frac{y}{\alpha})-f^{o}_{\mathcal{B}}(y)\geq 0$ for all $y\in[\alpha g^{p},\alpha\tau]$. Let $F(\cdot)$ denote the CDF of distribution $\mathcal{D}$. By definition of $f^{o}_{\mathcal{B}}(y)$ and $f^{p}_{\mathcal{B}}(y)$, we have
and by taking the derivative of $D_{\alpha}(\cdot)$, we immediately obtain that $D_{\alpha}(\cdot)$ is a weakly increasing function of $y$ and thus the condition only needs to hold on $y=\alpha g^{p}$. Observe that for $\alpha=g^{o}/g^{p}$ and $y=\alpha g^{p}=g^{o}$, we have $D_{\alpha}(y)=\frac{g^{o}}{g^{p}}\cdot f^{p}_{\mathcal{B}}(g^{p})-f^{o}_{\mathcal{B}}(g^{o})=g^{o}-g^{o}=0$ and thus we obtain that the opening action always achieves a $\frac{g^{o}}{g^{p}}$-local approximation. The proof is concluded by showing that
$\frac{g^{o}}{g^{p}}\leq\frac{c^{o}}{c^{p}}\cdot\big{(}1-\frac{c^{o}}{g^{p}}\big{)}.$
from which the claim follows.
Proof of 2.
By concavity, we have that $f^{p}_{\mathcal{B}}(y)\leq\alpha f^{p}_{\mathcal{B}}(\frac{y}{\alpha})$ for all $y\in\real$ and $\alpha\geq$ 1; thus, we only need to verify the condition for $y$ such that $f_{\mathcal{B}}(\frac{y}{\alpha})=f^{o}_{\mathcal{B}}(\frac{y}{\alpha})$; notice that this corresponds to $y\geq\alpha\tau>g^{p}$ and thus $f^{p}_{\mathcal{B}}(y)\leq y$. In other words, we only need to verify $f^{p}_{\mathcal{B}}(y)\leq\alpha f^{o}_{\mathcal{B}}(\frac{y}{\alpha})$ forall $y\geq\alpha\tau$.
by concavity of $f^{p}_{\mathcal{B}}(\cdot)$. Next, observe that
$h(\frac{c^{o}}{\alpha-1})\leq\frac{\alpha\cdot c^{o}-c^{p}}{\alpha-1}$
and for $\alpha=1+\frac{c^{p}}{c^{o}}$, this statement is equivalent to
$h(\frac{c^{o}\cdot c^{o}}{c^{p}})\leq\frac{c^{o}\cdot c^{o}}{c^{p}}$
which is clearly true, by definition of $h(\cdot)$.

### v2

@@ 5 Pandora’s Box with Partial Inspection (Minimization)
Formally, we consider an instance with $n$ partial inspection boxes (henceforth, PI-boxes) $\{\mathcal{B}_{i}\}_{i=1}^{n}$, where each box $\mathcal{B}_{i}=(\mathcal{D}_{i},c^{o}_{i},c^{p}_{i})$ is characterized by a distribution $\mathcal{D}_{i}$ over values; an opening cost $c^{o}_{i}>0$; and a peeking cost $c^{p}_{i}\in(0,c^{o}_{i}]$ [^10] Since we can only select an opened box, if $c^{p}_{i}\geq c^{o}_{i}$, then we can safely exclude the peeking action and the box reduces to a classical Pandora’s Box.. Each PI box $\mathcal{B}_{i}$ can be expressed as a Costly Information MDP with two possible actions, peeking and opening, where we can interpret the accept action after peeking as incurring a cost of $c^{o}$. An instance of PBPI corresponds to an instance of min-CICS over the corresponding MDPs. A pictorial representation of the different states of a box and the underlying MDP is shown in Figure 5.
Figure 5: In PBPI, a peeking box $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$ is initially closed. In order to learn the value realization realization $X\sim\mathcal{D}$, the decision maker can either peek into the box (at a cost of $c^{p}$) or open it (at a cost of $c^{o}$). To accept the box, the decision maker must first open it and then pay its (now known) value $X$.
A committing policy for PBPI needs to decide in advance (perhaps randomly) whether each box $\mathcal{B}_{i}$ will be directly opened or peeked into and then (potentially) opened. Our main result in this section is the following:
@@
The commitment gap of matroid-PBPI is at most $\sqrt{2}$.
In Section 5.1 we show a lower bound of $16/15$ for the commitment gap of matroid-PBPI (Example 2) and a trivial upper bound of $\varphi\approx 1.618$ (Lemma 5.2). We then prove Theorem 5.1 by showing that each PI box $\mathcal{B}$ admits a $\sqrt{2}\approx 1.414$ local approximation (Lemma 5.3) and employing Theorem 4.1; we note that the commitment achieving this guarantee decides whether each box will be directly opened or peeked before opened deterministically, based on the parameters of the box, and can be efficiently computed.
5.1 Lower and Upper Bounds on the Commitment Gap of PBPI
We begin by showing that that there are simple instances where all committing policies are sub-optimal up to a constant factor.
@@ 5.1 Lower and Upper Bounds on the Commitment Gap of PBPI
  • Since opening the second box is free, we can assume without loss that all policies start by opening it and observing its value, call it $y$. The optimal policy will simply open and accept the first box if $y=\infty$. But if $y=2$, it can peek into the first box and only open it if it contains a value of 00. The expected cost of this algorithm is $15/8$.
  • Now consider any policy that commits to opening or peeking into box $1$ before observing $y$. If it commits to opening the first box, it accepts the value of this box regardless of $y$, as box $1$ always has a value smaller than $y$. The expected cost of this policy is $2$. If it commits to peeking into the first box, then at $y=\infty$ it incurs a cost of $9/4$ due to having to pay the extra peeking cost; and at $y=2$ it incurs an expected cost of $7/4$ by opening the first box only if it contains a value of 00. It’s net cost is again $2$.
@@
Example 2 illustrates that the commitment gap of matroid-PBPI at least $(16/15)$. On the positive side, it is easy to obtain an upper bound of $2$. In particular, consider the policy that commits to peeking all boxes. This policy can mimic the optimal one as follows. Whenever the optimal algorithm peeks, so does this committing policy. Whenever the optimal algorithm opens without peeking, the committing policy peeks and then opens; on a box with costs $c^{p}_{i}$ and $c^{o}_{i}$, this policy pays $c^{p}_{i}+c^{o}_{i}$ or at most twice the amount $c^{o}_{i}$ paid by the optimal algorithm. In fact, we can further refine this argument by choosing the action we commit to more carefully: there always exists a simple commitment under which we can achieve a $\varphi\approx 1.618$ approximation to the optimal adaptive policy. The proof is presented in Appendix F.
@@
Consider any instance of matroid-PBPI and partition the $n$ boxes into two sets
@@
and $P=[n]\setminus O$. The policy that commits to directly opening the boxes in $O$ and peeking before opening the boxes in $P$ achieves a $\varphi$-approximation to the optimal (non-committing) policy.
We note that by using a global argument such as the one above, i.e. an argument where we charge each action of the committing policy directly to the optimal adaptive policy, one cannot improve on this upper bound of $\varphi$ (for example, consider a setting where all the boxes have $c^{p}=1$ and $c^{o}=\varphi$). In order to achieve a better guarantee, we would need to leverage our knowledge of the value distributions. In the next section, we achieve this by establishing local approximation guarantees for PBPI.
@@ 5.2 Local Approximation Guarantees for PBPI
Given a PI box $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$, let $g^{p}$ denote the water filling (a.k.a. Gittins) index of the policy that commits to peeking. Equivalently, $g^{p}$ is the solution to the equation $c^{p}=\operatorname{E}_{X\sim\mathcal{D}}\left[(g^{p}-X-c^{o})^{+}\right]$. We call $g^{p}$ the peeking index of the box.
@@
Let $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$ be a PI box with peeking index $g^{p}$. Let $\pi$ commit to the opening action whenever
@@
Definition 10 (Optimality Curves for PBPI).
The optimality curve of a PI box $\mathcal{B}=(\mathcal{D},c^{o},c^{p})$ for cost minimization is given by
@@
Figure 6: The optimality curves for a box $\mathcal{B}$ with opening cost $c^{o}=0.5$, value $X=0$ with probability $0.5$ and $X=2$ with probability $0.5$ and peeking cost $c^{p}=0.4$ (left) and $c^{p}=0.1$ (right). The optimality curve of $\mathcal{B}$ is given by the minimum of the three curves. Observe that the curves of the opening and peeking action intersect at a unique point $\tau$ and that unless $g^{p}<g^{o}$, the opening action dominates the peeking action.
Lemma 5.3 is a consequence of the following claims, establishing local approximation guarantees for the opening and peeking actions respectively.
@@
The proofs of the two claims are deferred to Appendix F. Note that once proven, they immediately imply an $\alpha$-local approximation with
@@
6 Additive Pandora’s Box (Minimization)
In this section, we study a different generalization of Pandora’s Box which we call Additive Pandora’s Box (henceforth, APB). In this setting, the value of each (additive) box is given by the sum of independent random variables, all of which need to be separately probed by paying a corresponding probing cost. The objective is to minimize the total sum of opening costs plus the additive value of the accepted box.
The primary motivation for studying APB is that it constitutes a special case of the notoriously challenging Pandora’s Shortest Path problem, where the edge weights of a given graph correspond to independent Pandora’s boxes and the objective is to accept a set of boxes that form a path between two given vertices $s,t$. Due to the fact that path constraints do not admit frugal algorithms, no results are currently known for this problem and it is left as an open direction by Singla (2017). It is not hard to see that APB corresponds to the special case of Pandora’s Shortest Path where the graph consists of parallel paths between $s$ and $t$; each path is then modeled via an additive Pandora’s Box (the components of which correspond to the path’s edges) and the path constraint translates to a single-selection constraint over the additive boxes.
Formally, an instance of APB consists of $n$-additive boxes $\{\mathcal{B}_{i}\}_{i=1}^{n}$. Each box $\mathcal{B}_{i}=(\vec{\mathcal{D}}_{i},\vec{c}_{i})$ is characterized by $k_{i}\geq 1$ random variables, each distributed independently according to $\mathcal{D}_{ij}$ and with an opening cost of $c_{ij}\geq 0$ for $j\in[k_{i}]$. The algorithm can observe the realization $V_{ij}\sim\mathcal{D}_{ij}$ by paying $c_{ij}$ and can accept the box $\mathcal{B}_{i}$ only after all the $k_{i}$ random variables have been observed, at a cost of $\sum_{j=1}^{k_{i}}V_{ij}$. At any point, the algorithm can adaptively select which additive box to explore, as well as which of its components to probe.
We can immediately cast this problem as a min-CICS instance where each additive box corresponds to an MDP $\mathcal{M}$ (which we call an additive-MDP), capturing all the different (adaptive) probing orders for the components of the box. A commitment for an additive-MDP corresponds to a protocol that determines which component to probe first, which second depending on the realization of the first value etc. Our main contribution is a to show that the commitment gap of APB is constant, and does not depend on the number of components.
Theorem 6.1.
The commitment gap of matroid-APB is at most $2$.
Our proof of Theorem 6.1 is existential and does not indicate how to efficiently construct the committing policy that achieves this bound. One could consider a restricted class of committing policies, which we call static commitment policies, where the components of each additive box are always probed in a fixed order. Trivial arguments suffice to show that static policies probing the components in increasing order of costs achieve a commitment gap of $k$. By considering the structure of water filling surrogate costs, we can obtain an improved bound on the gap:
Theorem 6.2.
Let $\mathcal{M}$ be any additive-MDP. For $k=2$, there exists a static commitment policy that achieves a $\varphi\approx 1.618$-pointwise approximation. For $k\geq 3$, the static policy of minimum index achieves a $\big{(}1+\lfloor\frac{k+1}{2}\rfloor\big{)}$-pointwise approximation.
In order to constructively achieve the approximation factor guaranteed by Theorem 6.2, one can compute the indices of the $k!$ static policies for each constituent MDP and commit to the one with the smallest index. We leave open the questions of whether it is possible to achieve the same approximation more efficiently, and whether one could efficiently compute a committing policy that admits a constant factor gap. The proof of Theorem 6.2 is presented in Appendix G.
Proof of Theorem 6.1
Our proof boils down to proving that each additive-MDP admits a $2$-pointwise approximation under a suitably chosen commitment and then employing 4.2 and Theorem 4.1. Our proof consists of two steps: first, we identify a simple structure for the surrogate costs under any commitment for APB and translate pointwise-approximation into a simple condition for the commitment. Then, we show that there exists some commitment satisfying this condition.
Step 1.
Let $\mathcal{M}$ be any additive-MDP with $k\geq 1$ components corresponding to values $V_{i}\sim\mathcal{D}_{i}$ and costs $c_{i}\geq 0$ for $i\in[k]$. Let $\pi\in C(\mathcal{M})$ be any commitment; observe that $\pi$ specifies which value to be probed first, which to be probed second depending on the first’s realization and so on until all $k$-values have been probed. Therefore, $\mathcal{M}^{\pi}$ is a tree-structured Markov chain of height $k$, the terminal states $t(\vec{V})$ of which correspond to all the possible realizations of the value vector $\vec{V}=(V_{1},\cdots,V_{k})$. Furthermore, the probability of reaching such a terminal state $t(\vec{V})$ is always equal to the probability of $\vec{V}$ being realized, independently of $\pi$’s probing protocol. Therefore, the water filling surrogate cost $W^{*}_{\mathcal{M}^{\pi}}$ for any commitment $\pi$ is obtained by sampling $\vec{V}\sim\vec{\mathcal{D}}$ and then returning a (commitment-specific) surrogate cost $\rho^{\pi}(t(\vec{V}))$. Due to this structure, the $\alpha$-pointwise approximation condition for additive-MDPs translates to proving that there exists some commitment $\pi\in C(\mathcal{M})$ such that
$\rho^{\pi}(t(\vec{V}))\leq\alpha\cdot\rho^{\pi^{\prime}}(t(\vec{V}))$
for all vectors $\vec{V}\sim\vec{\mathcal{D}}$ and all deterministic commitments $\pi^{\prime}\in C(\mathcal{M})$.
Up next, we will use the inductive definition of water filling amortization to obtain structure on these surrogate costs. Fix a commitment $\pi\in C(\mathcal{M})$ and the underlying Markov chain $\mathcal{M}^{\pi}$. For each state $s$ of $\mathcal{M}^{\pi}$, we define the following quantities:
  1. 1.
    The index $g^{\pi}(s)$ of state $s$ is defined as the water filling index of the sub-chain rooted at $s$, if all values that have been already probed have been realized to 00. In other words, $g^{\pi}(s)$ corresponds to the index of the committing policy that proceeds as $\pi$ after state $s$ is reached, without the additive cost of already observed values. Clearly, $g^{\pi}(t(\vec{V}))=0$ for all terminal states $t(\vec{V})$ and $g^{\pi}(r)=g^{\pi}$ for the root $r$ of $\mathcal{M}^{\pi}$.
  2. 2.
    The value $V^{\pi}(s)$ of state $s$ is defined as the sum of all the realized values in the trajectory from the root to $s$; clearly, we have that $V^{\pi}(r)=0$ and $V^{\pi}(t(\vec{V}))=\sum_{i=1}^{k}V_{i}$ for all terminal states $t_{\vec{V}}$.
The key observation is that when amortizing some state $s$, all the terminal states in it’s subtree will share the same additive offset $V^{\pi}(s)$, corresponding to the sum of realized values leading to $s$. It is immediate by the definition of the water filling amortization that such an additive offset does not affect the cost shares of vertex $s$. Also, recall that water filling amortization is performed in a bottom-up manner. Thus, the surrogate cost of a terminal state $t(\vec{V})$ immediately after one of its ancestors $s$ gets amortized will be precisely $V^{\pi}(s)+\max(g^{\pi}(s),\rho^{\prime}(t(\vec{V})))$, where $\rho^{\prime}(t(\vec{V}))$ is the surrogate cost of $t(\vec{v})$ immediately after it’s ancestor that is a child of $s$ was amortized. Performing this step for all ancestors of $t(\vec{V})$, we obtain that
$\rho^{\pi}(t(\vec{V}))=\max_{i=1}^{k+1}\bigg{(}V^{\pi}(s_{i})+g^{\pi}(s_{i})\bigg{)}$
where $(r=s_{1},s_{2},\cdots,s_{k},s_{k+1}=t(\vec{V}))$ is the unique path in $\mathcal{M}^{\pi}$ that reaches $t(\vec{V})$.
From this expression, it becomes clear that $\rho^{\pi}(t(\vec{V}))\geq\max(g^{\pi},\sum_{i=1}^{k}V_{i})$. Furthermore, if it is the case that $g^{\pi}\geq g^{\pi}(s)$ for all states $s$, then we can easily upper bound $\rho^{\pi}(t(\vec{V}))\leq g^{\pi}+\sum_{i=1}^{k}V_{i}$. We refer to commitments that have this property as root dominant. Combining everything, we then obtain that for any root dominant commitment $\pi$, any commitment $\pi^{\prime}$ and any terminal state $t(\vec{V})$, we have that
$\frac{\rho^{\pi}(t(\vec{V}))}{\rho^{\pi^{\prime}}(t(\vec{V}))}\leq\frac{g^{\pi}+\sum_{i=1}^{k}V_{i}}{\max(g^{\pi^{\prime}},\sum_{i=1}^{k}V_{i})}\leq 1+\frac{g^{\pi}}{g^{\pi^{\prime}}}.$
In other words, we have shown that any root dominant commitment $\pi$ will achieve an $\alpha$-pointwise approximation for $\alpha=1+g^{\pi}/\min_{\pi^{\prime}\in C(\mathcal{M})}g^{\pi^{\prime}}$. The proof of Theorem 6.1 is then completed by arguing that there exists a root dominant commitment $\pi$ that has minimum index across all committing policies.
Step 2.
Let $\pi$ be any commitment of minimum index. If $\pi$ is root dominant, then we are obviously done. Otherwise, there exists at least one state $s_{1}$ in $\mathcal{M}^{\pi}$ such that $g^{\pi}(s_{1})>g^{\pi}$. Among all these states, we consider one of minimum distance to the root. Notice that $s_{1}$ must have at least one sibling state $s_{2}$ with $g^{\pi}(s_{2})\neq g^{\pi}(s_{1})$; otherwise, the index of of the parent state of $s_{1}$ will be at least $g^{\pi}(s_{1})$ (by the inductive definition of indices) and thus $s_{1}$ won’t be a maximum height state satisfying the condition.
Let $\pi^{\prime}$ be the commitment that is obtained by substituting the sub-tree rooted at $s_{1}$ in $\mathcal{M}^{\pi}$ with the sub-tree rooted at $s_{2}$; notice that both of these subtrees correspond to commitments over the (same) set of unprobed values and thus such a substitution is always allowed (i.e. $\pi^{\prime}\in C(\mathcal{M})$ as well). Up next, we will show that $g^{\pi}\geq g^{\pi^{\prime}}$; then, $\pi^{\prime}$ will also be a minimum index commitment and we can apply the same process to it. Furthermore, each substitution reduces the heteromorphity of the underlying chain (i.e. the number of different child sub-trees that a state can have) so our iterative process is guaranteed to terminate after some finite number of steps, resulting in a commitment that is both root dominant and has minimum index.
We finally argue that $g^{\pi}\geq g^{\pi^{\prime}}$. Since $g^{\pi}$ is the water filling index of $\mathcal{M}^{\pi}$, it is by definition at least as large as the minimum surrogate cost of any valid amortization of $\mathcal{M}^{\pi}$. Let $b=\{b_{st}\}_{s,t}$ denote the water filling cost shares of $\mathcal{M}^{\pi}$ and $b^{\prime}=\{b^{\prime}_{st}\}_{s,t}$ denote the water filling cost shares of $\mathcal{M}^{\pi^{\prime}}$ Now, consider the following amortization $\hat{b}=\{\hat{b}_{st}\}_{s,t}$ of $\mathcal{M}^{\pi}$:
  • For each state $s$ that does not belong in the sub-tree rooted at $s_{1}$, let $\hat{b}_{st}=b^{\prime}_{st}$ for all terminals $t$.
  • For each state $s$ in the sub-tree rooted at $s_{1}$ (including $s_{1}$), let $\hat{b}_{st}=b_{st}$ for all terminals $t$.
Since all the states of $\mathcal{M}^{\pi}$ and $\mathcal{M}^{\pi^{\prime}}$ that are not ancestors of state $s_{1}$ have the same distribution over terminal states, this is indeed a valid amortization. Furthermore, any terminal state $t$ that is not a descendant of $s_{1}$ will receive precisely the same cost shares as it does in $\mathcal{M}^{\pi^{\prime}}$ and thus it’s surrogate cost will be at least $g^{\pi^{\prime}}$. On the other hand, every terminal state $t$ that is a descendant of $s_{1}$ receives the same cost shares as it did in $\mathcal{M}^{\pi}$ up until state $s_{1}$ gets amortized; by definition, this implies that it’s surrogate cost will be at least $g^{\pi}(s_{1})+V^{\pi}(s_{1})\geq g^{\pi}(s_{1})$. Thus, we conclude that $g^{\pi}\geq\min(g^{\pi^{\prime}},g^{\pi}(s_{1}))=g^{\pi^{\prime}}$ as desired.